OpenStudy (anonymous):

find the equation, in standard form of the line passing through the points (2,-3) (4,2)

OpenStudy (anonymous):

Do you know the formula for slope? and then the point-slope for of an equation of a line?

OpenStudy (anonymous):

point-slope form, not for

OpenStudy (anonymous):

isn't the slope form m=y1-y2/x2-x1?

OpenStudy (nincompoop):

first, solve for the slope then use that slope and one of the points into point-slope form solve for y then simplify into standard form

OpenStudy (anonymous):

yup! thats to find the slope. then you plug that answer into : y-y1=m(x-x1) you can use either point for the second equation

OpenStudy (anonymous):

and sorry, your slope one should be m=y2-y1/x2-x1

OpenStudy (anonymous):

what would you do with the y--3?

OpenStudy (anonymous):

the negatives cancel eachother out, making it positive